OpenStudy (anonymous):

A line has a slope of 2/3 and contains the points (x, 9) and (3, 7). The value of x is

OpenStudy (anonymous):

(3-x)=(7-9) do I distribute? or what do I do

OpenStudy (anonymous):

its 6 i figured it out

OpenStudy (deepolisnoob):

for this equation, you use your slope and the points you have. Your slope is 2/3 and your point is (3,7). What you need to do is plug in the values in a graph, chart, or just mentally. y=2/3x+b is the same as saying your y-value increases every time your x-value increases by 3. When your (3,7) turns to (x,9), your y-value went up by 2, meaning you add 3 to your x-value as well, making your second point (6,9)

OpenStudy (deepolisnoob):

ahh you already got it :P It's better when you figure it out on your own ;)

OpenStudy (anonymous):

ty